### v1

# Towards Vision-Language Geo-Foundation Model: A Survey

Abstract
Vision-Language Foundation Models (VLFMs) have made remarkable progress on various multimodal tasks, such as image captioning, image-text retrieval, visual question answering, and visual grounding. However, most methods rely on training with general image datasets, and the lack of geospatial data leads to poor performance on earth observation. Numerous geospatial image-text pair datasets and VLFMs fine-tuned on them have been proposed recently. These new approaches aim to leverage large-scale, multimodal geospatial data to build versatile intelligent models with diverse geo-perceptive capabilities, which we refer to as Vision-Language Geo-Foundation Models (VLGFMs). This paper thoroughly reviews VLGFMs, summarizing and analyzing recent developments in the field. In particular, we introduce the background and motivation behind the rise of VLGFMs, highlighting their unique research significance. Then, we systematically summarize the core technologies employed in VLGFMs, including data construction, model architectures, and applications of various multimodal geospatial tasks. Finally, we conclude with insights, issues, and discussions regarding future research directions. To the best of our knowledge, this is the first comprehensive literature review of VLGFMs. We keep tracing related works at https://github.com/zytx121/Awesome-VLGFM.
Index Terms:
Vision-language model, remote sensing foundation model, multimodal large language model
Figure 1: Timeline of representative VLGFMs. Purple entities indicate projects that currently play a significant role in advancing the development of VLGFMs. This field is experiencing rapid growth. For additional resources and daily updates, visit our GitHub page.
1 Introduction
Over the past decade, researchers have witnessed significant advancements in nearly all geospatial tasks, such as scene classification [1], object detection [2, 3], change detection [4], denoising [5], land use segmentation [6], disaster management [7], and geospatial localization [8], driven by deep learning and other artificial intelligence technologies. However, these models are specifically designed and trained for particular tasks, making them difficult to apply directly to other tasks. Even for the similar task, these models often exhibit poor generalization capabilities.
For example, remote sensing object detection is one of the core tasks for earth observation. It requires manually annotating the position and category of each object, which is a time-consuming and labour-intensive process. Remote sensing images (RSIs) are captured from an overhead perspective by spaceborne or airborne sensors, which present unique viewpoints compared to natural images, leading to the development of oriented object detection. Since this task uses rotated bounding boxes to represent objects, remote sensing datasets annotated with rotated bounding boxes, such as DOTA [9], are required to support its training. Furthermore, the model architecture [10], loss functions [11], post-processing functions, and acceleration operators [12] must be modified based on standard object detection [13]. From this perspective, the application scenarios of specific models seem quite limited, lacking the ability to generalize across different tasks, even between two similar tasks.
To reduce the resource wastage associated with training specific models from scratch for each task, foundational models [14] have emerged. These models are pre-trained on large-scale imagery, allowing them to handle various vision tasks with only fine-tuning small-scale custom datasets. In the field of remote sensing, prior studies on purely visual foundational models reveal immense potential for a generic model for earth observation, which is called Vision Geo-Foundation Models (VGFMs). VGFMs demonstrate remarkable generalization capabilities on a thorough evaluation, from single- to multi-modal, and static to temporal [15]. Although these models exhibit strong perceptual capabilities, they lack the ability to perform reasoning like humans. For instance, without corresponding annotated samples to support VGFM training, it cannot determine the specific function of a building in a remote sensing image by considering the surrounding environment and common sense, whereas humans can. Similarly, without annotated samples, VGFM cannot identify the brand or model of the car based on its features in aerial imagery, whereas humans can.
Recently, the evolution of Large Language Models (LLMs) has revolutionized human-machine interaction. LLMs like BERT [16] exploit vast quantities of textual data to develop reasoning skills, showing substantial generalization capabilities across diverse tasks in natural language processing. However, LLMs process only discrete textual data and cannot handle images, while visual foundation models, though capable of processing image data, lack inferential abilities. The concept of vision-language foundation models (VLFMs) has been introduced to bridge these gaps. These innovative models are designed to perform perception and reasoning, integrating inputs from text and images. Since the release of GPT-4 Vision, the research on VLFMs has been fervently pursued, inspired by its impressive capabilities. Substantial efforts have been made in VLFMs, which are typically divided into contrastive, conversational, and generative paradigms. Below, we will briefly introduce the most influential works in three directions. Specifically, CLIP [17] adopts the contrastive paradigm, projecting visual and textual information into a unified representational space, thereby facilitating a bridge for downstream vision-language tasks. LLaVA [18] embodies the conversational paradigm, endowing LLMs with the capability for context-aware dialogue across textual and visual modalities. Stable Diffusion [19], epitomizing the generative paradigm, leverages deep learning to produce high-quality, detailed images from textual descriptions, thereby advancing the capabilities in image synthesis and creative visual applications.
When VLFMs are applied to earth observation, this paper refers to them as Vision-Language Geo-Foundation Models (VLGFMs). To date, VLGFMs can also be categorized into contrastive, conversational, and generative types. Figure 1 lists the representative VLGFM developed and their release dates. It can be seen that VLGFMs first appeared in the second quarter of 2023. And now, the number of related works is in a period of rapid growth. Notably, current innovations in VLGFM are primarily focused on collecting training data, with relatively few modifications to the model architecture. Most efforts involve fine-tuning based on the LLaVA [18] and MiniGPT-4 [20] frameworks using custom-built remote sensing instruction-following sets.
As VLGFM continues to advance rapidly and demonstrate impressive results, tracking and comparing recent research on VLGFM is worthwhile. It transforms conventional human-computer interaction, which relies on predefined program interfaces, by enabling end-to-end interaction with humans through natural language dialogue. To the best of our knowledge, no comprehensive survey currently summarizes the latest developments in VLGFMs, including data pipelines, architectures, benchmarks, and capabilities. Our aim with this work is to fill this gap.
Contribution. In light of the rapid advancements and promising results of VLGFM, we have composed this survey to acquaint researchers with the fundamental concepts, principal methodologies, and current progress of VLGFMs. The survey extracts common technical details and covers the most representative works in the VLGFMs area. It also provides a comparative analysis of the background and related concepts, including VGFMs and remote sensing LLM-powered agents. Furthermore, we categorize the demonstrated capabilities of VLGFMs into three levels. To the best of our knowledge, we are the first survey on VLGFMs.
Survey pipeline. In Section 2, we provide background knowledge, including definitions, datasets, metrics, and related research areas. In Section 3, we conduct a comprehensive review of various methods based on different data collection approaches, network architectures, and capabilities. In Section 4, we identify challenges and future directions.
TABLE I: Model architectural of VLGFMs. Sort models by date of initial release. Abbreviations IS, IC, VC, IR, OC, REC, REG, RES, and GM in the table stand for image scene, image captioning, video captioning, image retrieval, object counting, referring expression comprehension, referring expression generation, referring expression segmentation, and geometric measurement, respectively. For each model, we indicate the LLM used in its best configuration as shown in the original paper ($\lozenge$: training from scratch; $\blacklozenge$: fine-tuning; $\blacktriangle$: fine-tuning with parameter-efficient fine-tuning techniques [21]; $\bigstar$: frozen).
Model LLM Visual Encoder Connector Capabilities Hardware
RemoteCLIP - ViT-L (CLIP) - IS, IR, 4*3090Ti (24G)
[22] OC
GeoRSCLIP - ViT-H (CLIP) - IS, IR 1*A100 (40G) +
[23] 1*A100 (80G)
RSGPT Vicuna-13B ViT-G (EVA) a Q-Former + VQA, IC 8*A100 (80G)
[24] a linear layer
GeoChat Vicuna-v1.5-7B ViT-L (CLIP) a two-layer MLP IS, VQA, IC, 3*A100 (40G)
[25] REC, REG
GRAFT - ViT-B (CLIP) - IS, IR, VQA, Unknown
[26] RES
SkyScript - ViT-L (CLIP) - IS, IR 4*A100 (80G)
[27]
SkyEyeGPT LLaMA2-7B ViT-G (EVA-CLIP) a linear layer IS, VQA, IC, 4*3090Ti (24G)
[28] REC, REG,
VC
EarthGPT LLaMA2-7B ViT-L (DINOv2) + a linear layer IS, VQA, IC, 16*A100 (80G)
[29] ConvNeXt-L (CLIP) REC, REG
LHRS-Bot LLaMA2-7B ViT-L (CLIP) a Vision Perceiver IS, VQA, IC, 8*V100 (32G)
[30] REC
RS-CapRet LLaMA2-7B ViT-L (CLIP) three linear layers IR, IC Unknown
[31]
H2RSVLM Vicuna-v1.5-7B ViT-L (CLIP) a two-layer MLP IS, IM, VQA, 16*A100 (80G)
[32] IC, REC, REG,
OC, GM, SP
RS-LLaVA Vicuna-v1.5-13B ViT-L (CLIP) a two-layer MLP VQA, IC 2*A6000 (48G)
[33]
SkySenseGPT Vicuna-v1.5-7B ViT-L (CLIP) a two-layer MLP VQA, IC, 4*A100 (40G)
[34] OVD, REC, REG
TABLE II: Evaluate data of VLGFMs. Abbreviations IC, VQA, VG, IS, and IR in the table stand for image captioning, visual question answering, visual grounding, image scene classification, and image retrieval, respectively (If more than 1 models use the dataset to evaluate a task, it is highlighted in blue. If more than 2 models use it, it is highlighted in red).
Model IC VQA VG IS IR Others
RemoteCLIP - - - PatternNet, EuroSAT, RemoteCount
[22] OPTIMAL-31, RSC11, RSITMD,
RSICB128, MLRSNet, RSICD,
RSI-CB256, RESISC45, UCM
WHU-earth, RS2800,
WHU-RS19
GeoRSCLIP - - - EuroSAT, AID, RSITMD, -
[22] RESISC45 RSICD
RSGPT UCM-Captions RSVQA-LR - - - RSIEval
[24] Sydney-Captions RSVQA-HR
GeoChat - RSVQA-LR - AID, - GeoChat-Bench
[25] RSVQA-HR UCM
GRAFT - RSVQA-HR - EuroSAT, BEN, EuroSAT, NAIP-OSM
[26] SAT-4, SAT-6 BEN
SkyCLIP - - - PatternNet, EuroSAT, RSITMD, SkyScript
[27] AID, fMoW, RESISC45, RSICD,
MillionAID, RSI-CB256 UCM
SkyEyeGPT UCM-Captions, RSVQA-LR, RSVG, - - -
[28] Sydney-Captions, RSVQA-HR DIOR-RSVG
RSICD, CapERA
EarthGPT NWPU-Captions CRSVQA DIOR-RSVG RESISC45 - -
[29] RSVQA-HR MAR20 CLRS, NaSC-TG2
LHRS-Bot - RSVQA-LR, RSVG, AID, WHU-RS19, - -
[30] RSVQA-HR DIOR-RSVG SIRI-WHU, EuroSAT,
RESISC45, fMoW,
METER-ML
RS-CapRet UCM-Captions, - - - RSICD -
[31] Sydney-Captions, UCM
NWPU-Captions,
RSICD
H2RSVLM - RSVQA-LR DIOR-RSVG NWPU, METER-ML, - RSSA
[32] RSVQA-HR SIRI-WHU, AID, HqDC-Instruct
WHU-RS19
RS-LLaVA UCM-Captions, RSVQA-LR, - - - -
[33] UAV RSIVQA
SkySenseGPT RSVQA-LR AID, WHU-RS19, FIT-RSFG
[34] RSVQA-HR SIRI-WHU FIT-RSRC
2 Background
Overview. In this section, we first introduce the conceptual definition of VLGFM and compare it with related concepts. Then, we provide a historical overview of VLGFM and highlight several representative models. Next, we present the commonly used benchmark datasets and metrics in the VLGFM field. Finally, we review related research domains.
2.1 Concept Definition
Foundation models is a kind of model with broad applicability and general capabilities after training on large-scale data. These models typically have a large number of parameters and strong learning abilities, enabling them to excel in a variety of tasks. Although foundation models may not be ubiquitous at present, they appear poised to become the basis of widespread technological innovations and exhibit the key hallmarks of a general-purpose technology [35]. The field surveyed in this paper can handle multiple tasks simultaneously, and vision-language models that support only a single task are outside the scope of our research. Please refer to [14] for those interested in task-specific vision-language models.
Geo-foundation models are a class of models specifically designed for processing geospatial data through visual information. These models utilize various types of geospatial visual data, such as remote sensing imagery, satellite photos, and aerial images, to perform detailed analysis and support a range of geographic applications. Although this research is currently a popular topic, it is not the focus of this paper. For those interested in purely visual geo-foundation models, please refer to [36] for further information.
Vision-language geo-foundation models are a specialized subset of artificial intelligence models designed for processing and analyzing geospatial data by integrating visual and linguistic information. These models can handle diverse geospatial data sources, such as remote sensing imagery, geographic information system data, and geo-tagged text, leveraging their cross-modal processing capabilities to understand and integrate different types of geospatial information. By combining visual and linguistic modalities, VLGFMs can perform more comprehensive and accurate analyses of geospatial data, making them highly valuable for complex earth observation tasks. This paper focuses on these VLGFMs, excluding those limited to the visual modality or specific tasks. As far as we know, this is the first exhaustive literature review of VLGFMs.
2.2 History and Roadmap
Before introducing the VLGFM details, it is necessary to review its progress. In Figure 1, we summarize the timeline of VLGFM and list some representative works. The development of VLFMs in the remote sensing domain started relatively late. It was not until the emergence of the works LLaVA [18] and MiniGPT-4 [20] in 2023 that VLGFM-related work began to take shape. Most existing conversational VLGFMs have been implemented based on the open-source frameworks released by these two works. Since June 2023, the first batch of VLGFMs began to emerge. Among them, RemoteCLIP [22] was the first contrastive VLGFM, supporting image scene classification and image-text retrieval tasks. RSGPT [24] was the first conversational VLGFM, supporting image captioning and visual question answering tasks. RS5M [23] was the first million-scale remote sensing image-text pair dataset made publicly available. Since October 2023, the number of VLGFM works has steadily increased. Among them, DiffusionSat [37] is the first high-quality text-to-image generation work in the remote sensing field. It can control the generation of RSIs using conditions such as latitude and longitude, season, cloud cover, and ground sampling distance (GSD). GeoChat [25] is the first conversational VLGFM to support visual grounding task., while SkyEyeGPT [28] is the first to support remote sensing video captioning task. H2RSVLM [32] is the first to introduce the concept of trustworthiness, where the model refuses to answer uncertain questions. SkySenseGPT [34] introduces the concept of graphs to enhance VLGFM’s ability to perceive and reason relationships between remote sensing targets as well as its capability for complex comprehension. It is worth mentioning that research on VLGFMs is more akin to a data-centric project rather than a model-centric one. As seen from Table I, the structure of each model is quite similar, with no significant modifications. However, Tables II shows that the data selection for different VLGFMs varies considerably.
2.3 Tasks, Datasets, and Metrics
$\bullet$ Tasks. VLGFMs are capable of addressing a wide range of earth observation tasks, including image scene classification (IS) [38], image retrieval (IR) [39], image captioning (IC) [40], visual question answering (VQA) [41], and visual grounding (VG) [42]. Among these tasks, the VQA task is particularly special as it serves as a means to evaluate most multi-modal tasks in VLGFMs. Specifically, the inputs of VQA include images and text prompts, and the output is text. By using different text prompts, answers for various tasks can be obtained. VG is a more practical task in real-world applications and is mainly divided into two sub-tasks: referring expression comprehension and referring expression generation. Detailed introduction to the various capabilities of VLGFMs will be provided in Section 3.3.
$\bullet$ Datasets. In Table II, we list the datasets commonly used in experimental validation of existing VLGFMs and categorize them into three types based on usage frequency: recommended, common, and rare. Red, blue, and black fonts represent these categories. For the image captioning task, the recommended datasets are UCM-Captions and Sydney-Captions [40]; the commonly used datasets are RSICD [43] and NWPU-Captions [44]; and the rare datasets are CapERA [45] and UAV [46]. For the VQA task, the recommended datasets are RSVQA-LR and RSVQA-HR [41]. The recommended dataset for the visual grounding task is DIOR-RSVG [47]; the commonly used dataset is RSVG [42]; and the rare dataset is MAR20 [48]. For the image scene task, the recommended datasets are EuroSAT [49], RESISC45 [50], WHU-RS19 [51], SIRI-WHU[52], and AID [53]; the commonly used datasets are PatternNet [54], RSI-CB256 [55], METER-ML [56] and fMoW [57]. For the image retrieval task, the recommended datasets are RSITMD [39], RSICD [43], and UCM [58]. At the same time, some VLGFMs have also released their own evaluation sets constructed for different specific capabilities.
$\bullet$ Metrics. For the image captioning task, the commonly used metrics include BiLingual Evaluation Understudy (BLEU), Recall-Oriented Understudy for Gisting Evaluation (ROUGE-L), Metric for Evaluation of Translation with Explicit ORdering (METEOR), and Consensus-based Image Description Evaluation (CIDEr). In VQA and image scene tasks, the primary metric used is Accuracy. For the visual grounding task, the commonly used metric is Accuracy@0.5, which considers a prediction accurate if the predicted bounding box overlaps with the ground-truth box by more than 0.5 Intersection over Union (IoU). The commonly used metric for the image retrieval task is the recall of top-K (R@K), which measures the ratio of queries that successfully retrieve the ground truth within the top K.
Additionally, experts can be employed to score the outcomes of responses. For instance, RSGPT [24] utilizes experts to categorize the results of captioning tasks into four distinct levels. Since the advent of large language models, efforts have been made to leverage these models for scoring the outcomes of various vision-language multimodal tasks. For choice questions, the output of some models may not always meet the requirements due to imperfect instruction-following capabilities. Their responses may be completely unrelated to the questions, often failing to convey the intended meaning. Their responses may be completely unrelated to the questions, consisting of meaningless numbers and special symbols. MMBench [59] employs ChatGPT [60] as a judge to identify the option most similar to the model’s output and regard it as the model’s prediction result.
2.4 Related Research Domains
$\bullet$ Vision geo-foundation models. VGFMs are also called remote sensing pre-trained models. The primary difference between these models and VLGFM is that their input is purely visual, without any modeling of textual information. RVSA [61] trained a conventional vision transformer with 100 million parameters on RGB RSIs and developed a novel rotational resizing window attention mechanism to fine-tune the model for downstream tasks. Unlike MAE-based methods that infer the entire image from a few visible patches, RingMo [62] implemented a masked image modelling strategy in their pre-trained model, considering all image patches, whether masked or unmasked. The success of these preliminary studies demonstrates the significant potential of pre-trained models in remote sensing applications. Recently, foundational visual models for spectral [63], multimodal [15], and synthetic aperture radar (SAR) [64] images have been proposed, attracting significant attention in the VGFM field.
$\bullet$ Remote sensing LLM-powered agent. The concept of an LLM-powered agent emphasizes leveraging the capabilities of LLMs to decompose tasks into multiple steps and then accomplish specific tasks by invoking appropriate tools. The most well-known agent framework currently is LangChain [^1] https://www.langchain.com, which allows developers to integrate large language models with external computation and data sources [65, 66, 67, 68]. In the field of remote sensing interpretation, an LLM-powered agent [69] would utilize pre-prepared toolkits, including object detection and classification algorithms, to complete visual-language multimodal tasks. Recently, RS-Agent [70] treats conversational VLGFMs, such as GeoChat and LHRS-Bots, as tools, utilizing them to achieve RSI captioning functionality.
3 Methods: A survey
Overview. In this section, we systematically explore how data is gathered and utilized to train these models (Section 3.1), examine the architectural choices and modifications made to enhance their performance (Section 3.2), and precisely define their capabilities by the variety of tasks they can handle (Section 3.3). It is important to note that our study highlights and compares only the most representative and influential works in the field. This approach ensures a comprehensive understanding of the current state of VLGFM development while emphasizing the key advancements and methodologies that have shaped their evolution.
TABLE III: Training Data of VLGFMs. Abbreviations IC, VQA, IS, Det., Seg., and VG in the table stand for image captioning, visual question answering, image scene classification, object detection, segmentation and visual grounding, respectively (If more than 1 models use the dataset to evaluate a task, it is highlighted in blue. If more than 2 models use it, it is highlighted in red).
Model IC VQA IS Det./Seg. VG Others
RemoteCLIP RSICD, RSITMD, - - DOTA, DIOR, HRRSD, RSOD, - -
[22] UCM-Captions LEVIR, HRSC, VisDrone,
AU-AIR, S-Drone, CAPRK,
Vaihingen, Potsdam, iSAID, LoveDA
GeoRSCLIP - - fMoW, BEN, - - RS5M
[23] MillionAID
RSGPT - - - - - RSICap
[24]
GeoChat - FloodNet, RESISC45 DOTA, DIOR, FAR1M, - -
[25] RSVQA-LR SAMRS
GRAFT - - - - - NAIP,
[26] Sentinel-2
SkyCLIP - - - - - SkyScript
[27]
SkyEyeGPT RSICD, RSITMD, ERA-VQA, - - RSVG, DOTA-Conversa
[28] UCM-Captions, RSVQA-LR, DIOR-RSVG, DIOR-Conversa
Sydney-Captions, RSVQA-HR, RSPG UCM-Conversa
NWPU-Captions, RSIVQA Sydney-Conversa
CapERA
EarthGPT RSICD, RSITMD, FloodNet, DSCR, EuroSAT, DOTA, DIOR, FAR1M, NWPUVHR10, DIOR-RSVG MMRS-1M
[29] UCM-Captions, RSVQA-LR, UCM, WHU-RS19, HRRSD, RSOD, UCAS-AOD, VisDrone,
Sydney-Captions, RSVQA-HR, RSSCN7, RESISC45, AIR-SARShip-2.0, SSDD, HRISD, HIT-UAV,
NWPU-Captions CRSVQA FGSCR-42 Sea-shipping, Infrared-security, Aerial-mancar,
HIT-UAV, Double-light-vehicle, Oceanic ship
LHRS-Bot RSICD,RSITMD, RSVQA-LR RSITMD, NWPU, - RSVG LHRS-Align,
[30] UCM-Caption RSVQA-HR METER-ML,UCM, DIOR-RSVG LHRS-Instruct
NWPU-Captions fMoW
RS-CapRet RSICD, - - - - -
[31] UCM-Captions,
Sydney-Captions,
NWPU-Captions
H2RSVLM - RSVQA-LR fMoW, RSITMD, DOTA, FAIR1M, LoveDA, DIOR-RSVG CVUSA, CVACT,
[32] MillionAID, NWPU, MSAR, GID, FBP, BANDON,
METER-ML, UCM DeepGlobe, CrowdAI MtS-WH
RS-LLaVA UCM-Captions, RSVQA-LR, UCM - - -
[33] UAV RSIVQA -
SkySenseGPT FloodNet RSITMD, DOTA RSG
[34] RSVQA-LR UCM, DIOR
Earth-VQA NWPU FAIR1M
TABLE IV: Data Pipeline of VLGFMs.
Model Data pipeline
RemoteCLIP Retrieval Data: use it directly. Detection Data: generate 5 distinct captions that describe the objects in the image
[22] (the first 2 captions are generated according to the target location, and the remaining 3 captions are generated by
considering the number of different object categories). Segmentation Data: first convert it to boxes, then use the
same pipeline as the detection data.
RS5M PUB11: utilize regex to identify image-text pairs containing the RS-related keywords, utilize fastdup for invalid
[23] image checking and deduplication. Then, clean the dataset using the CLIP score and the RS image classifier. RS3:
generate 20 candidate captions per image using tuned BLIP2 and rank the top 10 results using CLIP ViT-H/14.
Then, re-rank these top 10 results using CLIP Resnet50x64 to obtain the top 5 captions.
RSGPT RSICap: 5 RS experts annotate 2,500 images from the detection dataset, focusing on (1) image attributes like
[24] type, and resolution; (2) object details such as quantity, color, shape, size, and position; (3) annotate the overall
scene before specific objects. RSIEval: For Cap. task, same with RSICap (1 caption per image). For VQA task,
product QA pairs for each image (9 questions per image), categorized into object-related, image-related, scene-
related, and reasoning-related questions.
GeoChat Attribute extraction: Objects’ categories are sourced from the SAMRS dataset. Objects’ colors are obtained by
[25] K-Means. Objects’ relative sizes are small, normal, or large. Objects’ relative positions are assessed using a 3x3
grid overlay on images, and the relationship of objects’ bounding boxes determines their relations. Expression
Generation: employ predefined textual templates based on objects’ attributes to emulate natural language
expressions. Instruction Generation: generated referring expressions are input into an LLM, which outputs 3
kinds of question-answer pairs, i.e. grounding image description, referring expression, and region captioning.
GRAFT Ground Images: We sourced outdoor ground images with precise geo-tags from diverse regions on Flickr,
[26] ensuring a uniform location distribution avoiding duplicates, and filtering out indoor photos with a classifier.
Satellite Images: Satellite images are sampled with centres at the geotags of ground images, with each assigned
all non-duplicate ground images within its bounds. Overlaps between satellite images are minimized (at least 112
pixels apart). For any location with more than 25 ground images, a random subset of 25 is selected.
SkyScript Images are from Google Earth Engine (GEE), and the geographic features are gathered from OpenStreetMap
[27] (OSM). Connect images with appropriate OSM semantics using a two-stage tag classification approach with
CLIP tag embeddings. Converting tags into a caption by first connecting the key and the value with a connecting
word (e.g., ”of”, ”is”) and then connecting multiple tags with commas or ”and”. Filtering out uncorrelated image
-text pairs by CLIP model to derive embeddings and compute cosine similarity.
SkyEyeGPT Single-task Image-text Instruction: integrate 5 image captioning, 1 video captioning, and 4 VQA datasets.
[28] Following the miniGPT-v2, create an phrase grounding dataset. Multi-task Conversation Instruction: mix the
corresponding captioning and VQA datasets to get UCM-Conversa and Sydney-Conversa instruction. Using the
DIOR-RSVG and DIOR to construct DIOR-Conversa instruction which contains visual grounding, phrase
grounding, and referring expression generation tasks. Leverage RSIVQA and the DOTA to build a conversation
instruction, DOTA-Conversa, that includes VQA and phrase grounding tasks. Data is manually verified and
selected by humans to ensure the high quality of the instruction.
EarthGPT Convert 10 classification datasets, 5 image captioning datasets, 4 VQA datasets, 17 object detection datasets, and
[29] 1 visual grounding dataset into instructions using prompt templates.
LHRS-Bot LHRS-Align: images are from GEE, and the geographic features are gathered from OSM. Filtering out irrelevant
[30] keys and balance semantic information. Using Vicuna-v1.5 to generate image captions by summarising the key
-value tags of the corresponding images. LHRS-Instruct: For public RS caption dataset, a rigorous data cleaning
process is implemented, including deduplication, caption length computation, and relevance assessment via CLIP
score. Several in-context examples with the five captions per image prompt Vicuna-v1.5 to generate object-focused
conversations. The most relevant conversations are filtered using predefined criteria and expert evaluations. For
LHRS-Align, selecting the attribute-rich RS images from the LHRS-Align dataset, bounding boxes for objects
within these images are calculated using OSM database spatial coordinates. This enriched data prompts GPT-v4
to create complex instruction data featuring visual reasoning, detailed descriptions, and object location and count
conversations. The final step involves a manual cleaning of the generated results.
H2RSVLM HqDC-1.4M: designed appropriate prompts to guide the Gemini-Vision model to provide detailed descriptions of
[32] remote sensing images, including the type and resolution of the images, as well as information on the attributes,
quantities, colors, shapes, and spatial positions of objects in the images. HqDC-Instruct: combined the
aforementioned descriptions with the original object detection annotations of the dataset and input them into the
language-only Gemini. Used prompts and few-shot examples to guide Gemini in generating multi-turn conversation
and reasoning data. RSSA: Designed question-answer pairs based on two RS object detection datasets, some of
which questions cannot be answered using the information in the images alone. RS-Specialized-Instruct: A remote
sensing-specific instruction-following dataset created using some existing remote sensing datasets. RS-ClsQaGrd-
Instruct: a multi-task instruct dataset constructed by public remote sensing datasets, covering tasks such as scene
classification, VQA, and visual grounding.
RS-LLaVA Convert 2 image captioning datasets and 2 VQA datasets into instructions using prompt templates.
[33]
SkySenseGPT FIT-RS: enhance a scene graph generation remote sensing dataset, which contains complex relationships between
[34] remote sensing objects.
Motivation of survey organization. Firstly, there is currently no comprehensive review focused on VLGFM. Secondly, we believe that the current emphasis in VLGFM research should be more on data rather than fundamentally altering the architecture of general large models and training them from scratch, which may ultimately prove to be futile. There are several reasons. (1) The available image-text data in remote sensing is significantly less than in computer vision. This gap makes it infeasible to support the training of a remote sensing image-text alignment model from scratch. Therefore, constructing high-quality, large-scale remote sensing image-text datasets remains a top priority for VLGFM research in the foreseeable future. (2) Training VLGFMs from scratch requires substantial time and resources. Without significant resource investment, any short-term lead is likely to be quickly overtaken by the rapidly evolving general-purpose multimodal models. (3) VLGFMs represent the application of large general models in the remote sensing domain, falling under the category of specialized artificial intelligence. The key is to leverage the extensive prior knowledge in the remote sensing field to address specific, real-world problems. Thus, fine-tuning networks using high-quality remote sensing image-text datasets to enable the model to generalize and apply knowledge effectively is a practical approach. In summary, our investigation focuses on providing a more comprehensive review and objectively analyzing the technical approaches of existing VLGFM works.
3.1 Data Pipeline
VLGFMs require finetune with remote sensing image-text pairs. The collection of training samples is primarily divided into two categories. One approach involves downloading raw RSIs and obtaining their corresponding annotations through various websites or general multimodal models. The other approach leverages existing remote sensing dataset resources by converting the original annotations into textual captions using prompt templates. In the following sections, we will introduce these two methods in detail.
$\bullet$ Data collection from scratch. As shown in Table III, three VLGFMs have built their training datasets from scratch rather than utilizing existing ones. RSGPT [24] prioritizes the quality of training samples over their quantity. Hence, it employs five experts to annotate images, producing high-quality captions manually. As a result, the model achieves satisfactory performance with fine-tuning on only 2,500 remote sensing image-text pairs. GRAFT [26] ingeniously utilizes the geo-tag information of images to link ground-level image captions collected from social media with RSIs. This approach enables the acquisition of many remote sensing image-text pairs without the need for manual annotation of RSI captions. SkyScript [27] utilizes a well-known open-source collaborative project, OpenStreetMap [^2] https://www.openstreetmap.org (OSM), to construct image captions. OSM allows users to contribute and edit geospatial data, such as road networks, trails, cafes, railway stations, etc. It provides a comprehensive and freely accessible world map and detailed user annotations. SkyScript filters these annotations and combines various tags to generate high-quality image captions. In LHRS-Bo [30], the LHRS-Align also acquires tags of geographic features from OSM and generates image captions based on these tags using language-only LLM, Vicuna-v1.5. More details can be found in Table IV.
Figure 2: Architecture of representative VLGFMs: (a) Contrastive, (b) Conversational, (c) Generative.
$\bullet$ Data enhancement using existing datasets. Most works choose to modify existing remote sensing datasets, which is significantly less labour-intensive compared to building datasets from scratch. Consequently, researchers can focus more on improving the model’s structures. RemoteCLIP [22] was the first to propose using text templates to convert the most common object detection annotations in the field of remote sensing interpretation into image captions, generating a large number of remote sensing image-text pairs suitable for training large models. Subsequently, EarthGPT [29] also adopted the same approach. RS5M [23], which currently contains the largest image-text pairs, includes 5 million RSIs. It employs the general multimodal model BLIP2 [71] to generate image captions and then uses CLIP [72] to select the top 5 highest-scoring captions. GeoChat [25] constructs a remote sensing image-text pair dataset based on the segmentation dataset generated using SAM [73] technology. It first uses the target’s color, size, and location to build basic expressions and then inputs these expressions into a language-only LLM to generate question-answer pairs. SkyEyeGPT [28] constructs a multi-task conversation instruction dataset by combining an object detection dataset with a VQA dataset built using the object detection data. This approach supports training for both VQA and visual grounding tasks. For more details, please refer to Table IV.
3.2 Architecture
VLGFMs can be briefly classified into three categories based on their input-output mechanisms. Contrastive VLGFMs [22, 23, 26, 27] take both text and images as inputs and produce a similarity measure between them. This similarity is crucial for image-text retrieval and zero-shot scene classification applications. Conversational VLGFMs [24, 25, 28, 29, 30, 31, 32] also process text and images as inputs but output textual responses, leveraging the powerful language capabilities of LLMs to support tasks such as captioning and visual question answering. Additionally, dense prediction tasks can be supported by appending task-specific heads to the LLM. On the other hand, generative VLGFMs [37, 74] are conditioned on either text or images, producing images as outputs. These models utilize conditional diffusion processes to generate controlled synthetic RSIs. This section will provide a detailed exposition of the model architectures for these three types of VLGFMs.
3.2.1 Contrastive VLGFMs
A typical contrastive VLGFM comprises two main components: an image encoder and a text encoder. A schematic of this architecture is depicted in Figure 2. Image encoder uses a convolutional neural network or a vision transformer to encode images into a feature vector. Parallel to the vision encoder, the text encoder (typically a BERT model) processes textual descriptions into another feature vector. The core idea behind contrastive VLGFM is to align the image and text embeddings in a shared multidimensional space. During training, contrastive VLGFM uses a contrastive learning objective, specifically a noise contrastive estimation technique variant. The model is trained to maximize the similarity between positive image-text pairs and minimize the similarity between negative pairs. The large-scale training allows contrastive VLGFM to understand various visual concepts described in natural language robustly. CLIP [72] is one of the pioneering models for vision-language pre-training. It employs the InfoNCE loss to align text and images in the feature space. Benefiting from a dataset comprising 400 million image-text pairs, CLIP has achieved significant success in various multimodal downstream tasks. BLIP [75] is the first multimodal model that simultaneously addresses both understanding and generation of image-text content, which is jointly pre-trained on three vision-language tasks: image-text contrastive learning, image-text matching, and image-conditioned language modelling.
In the domain of earth observation, some outstanding vision-language pre-training models have also emerged. RemoteCLIP [22] is the first true contrastive VLGFM. It utilizes existing remote sensing retrieval, detection, and segmentation datasets to construct image-text sample pairs, supporting model training. GeoRSCLIP is a vision-language pre-training model released alongside the RS5M dataset [23]. The authors did not provide extensive details about it in the paper. GRAFT [26] is a significant attempt to address the issue of insufficient samples in remote sensing image-text data. It utilizes a vast collection of ground-level images from Flickr [^3] https://www.flickr.com as a bridge, pairing RSIs that correspond to the geographic coordinates of these ground images with the titles of the ground images. This approach allows for the construction of a large-scale dataset of remote sensing image-text pairs without the need for additional manual annotation. SkyScript [27] utilizes the ground sample distance (GSD) of OSM data to annotate images appropriately. For instance, house numbers cannot be discerned from RSIs regardless of their resolution, and utility poles are not visible in RSIs with a GSD of 10. However, these contrastive VLGFMs are primarily data-centric efforts, with no significant modifications made to the model structures.
3.2.2 Conversational VLGFMs
A typical conversational VLGFM comprises three main components: a pre-trained visual encoder, a pre-trained LLM, and a modality interface that connects them. Drawing an analogy with human sensory and cognitive processes, the visual encoder functions similarly to the human eye, receiving and preprocessing optical signals. The LLM, akin to the human brain, can understand these processed signals and perform reasoning tasks. The interface between them serves to align the visual and linguistic modalities. A schematic of this architecture is depicted in Figure 2. In this section, we will describe each of these components in detail.
$\bullet$ Visual encoder. The visual encoder compresses images into more compact representations. Training this network from scratch is time-consuming; therefore, a common approach is to utilize the image encoder from an already text-aligned contrastive VLGFM. Aligning the visual encoder with textual semantics is advantageous for subsequent integration with LLMs.
Table I presents the visual encoders commonly used in current VLGFM implementations. Besides the standard CLIP image encoder [72], some studies have explored other variants. For instance, RSGPT and SkyEyeGPT employ the EVA-CLIP encoder [76], which is trained with improved training techniques. Additionally, EarthGPT utilizes two encoders simultaneously and introduces the ConvNext encoder [77], a convolution-based architecture designed to exploit higher resolutions and multi-level features.
$\bullet$ Pre-trained LLM. Training an LLM from scratch entails prohibitive computational costs for most individuals, making it more effective and practical to begin with a pre-trained one. Through pre-training on vast amounts of data collected from various places, the LLM acquires rich world knowledge and demonstrates strong capabilities in generalization and reasoning. Table I summarizes commonly used and publicly available LLMs, all of which fall under the category of causal decoders [60]. The LLAMA series [18] and the Vicuna family [78] represent some of the most notable open-source LLMs that have garnered considerable attention within the academic community. It is important to note that these LLMs were primarily pre-trained on English corpora, which limits their multilingual capabilities, such as for Chinese. Additionally, increasing the parameter size of LLMs also brings additional benefits, akin to increasing input resolution. Specifically, RSGPT [24] found that simply expanding the Vicuna from 7B to 13B parameters can significantly enhance performance across various benchmarks.
$\bullet$ Vision-language connector. Since LLMs are limited to text perception, bridging the gap between natural language and visual modalities is necessary. However, training large multimodal models in an end-to-end manner would be costly. A more practical approach is introducing learnable vision-language connectors between pre-trained visual encoders and LLMs. The module efficiently projects image information into a space that LLMs can understand. Based on the fusion method between visual and language information, these connectors can generally be divided into token-level and feature-level connectors.
For token-level connectors, the output of the encoder is transformed into tokens and concatenated with text tokens before being fed into the LLM. A common and feasible solution is to use a set of learnable query tokens to extract information in a query-based manner, initially implemented with BLIP-2 [71] and subsequently adopted by various works on VLGFMs [24]. This query-based approach compresses visual tokens into a smaller number of representation vectors. In contrast, some methods employ multi-layer perception (MLP) to bridge the modal gap. For instance, the LLaVA series only use linear layers [28, 29, 31] or two-layer MLPs [25, 32] to project visual tokens and align feature dimensions with word embeddings.
For feature-level connectors, it is required to incorporate additional modules to facilitate deep interaction and integration between textual features and visual features. For instance, Flamingo [79] introduces additional cross-attention layers between the frozen transformer layers of the LLM, thereby utilizing external visual cues to enhance language features. Similarly, LHRS-Bot [30] attempts to introduce a set of learnable queries for each level of image features. These queries aim to encapsulate the semantic information of each level of visual features through a series of stacked cross-attention and MLP layers.
$\bullet$ Pre-training. Pre-training primarily aims to align different modalities and learn multimodal world knowledge in the initial training phase. The pre-training phase typically requires large-scale image-text pair data, such as caption data. For a given image, the model is trained to autoregressively predict the image’s caption using a standard cross-entropy loss. A common approach during pre-training is to keep pre-trained modules (such as the visual encoder and LLM) frozen and train the vision-language connector [24]. The idea is to align the visual and linguistic modalities without losing the knowledge acquired during pre-training. Some methods [28, 29, 30] also unfreeze LLM to allow more trainable parameters for alignment. ShareGPT4V [80] found that having high-quality caption data during pre-training and unlocking visual encoding capabilities can promote better alignment. In the existing VLGFM landscape, only H2RSVLM [32] has attempted to unfreeze all modules simultaneously for pre-training.
$\bullet$ Instruction-tuning. Instruction-tuning aims to teach models to understand user commands better and complete desired tasks. Through this approach, LLMs can be generalized to support unseen tasks, enhancing zero-shot performance. This simple yet effective idea has spurred the success of subsequent NLP works, such as InstructGPT [81]. Supervised fine-tuning methods typically require much training data to train task-specific models. Instruction-tuning methods reduce the reliance on large-scale data and can accomplish specific tasks through prompt engineering. Due to the flexible format of instructional data and the variability in templates across different tasks, collecting data samples is often more challenging and costly. Recent studies have shown that the quality of instructional fine-tuning samples is more critical than quantity. [82] discovered that high-quality instruction-tuning data can yield better performance, even if the quantity of samples is not the largest.
3.2.3 Generative VLGFMs
A typical generative VLGFM comprises three main components: a text encoder, a pre-trained conditional diffusion model, and an image decoder. A schematic of this architecture is depicted in Figure 2. The working mechanism of diffusion models consists of two phases. Initially, noise is introduced into the dataset, which is a crucial step in the forward diffusion process, followed by a systematic reversal of this process. In other words, the method involves identifying a stochastic differential equation (SDE) [83] whose process converges to a standard normal distribution. This equation transforms the data, mapping it onto the standard normal distribution. Subsequently, samples that conform to the original data distribution are generated by sampling from this distribution and applying the inverse of the SDE to the sampled data.
Conditional diffusion models are a class of generative models that extend the fundamental principles of diffusion processes to incorporate conditional information, thereby enabling the generation of data that adheres to specific conditions or constraints. In the field of remote sensing, DiffusionSat[37] successfully enabled the model to generate satellite images of specific locations and seasons by incorporating information such as latitude, longitude, and the time of capture during the training process of the diffusion model. Furthermore, CRS-Diff [74] attempted to guide the diffusion model in generating RSIs with similar layouts by using binary masks or road sketches, which can be used to generate samples that are difficult to collect in the real world. Due to limitations in the number of training samples and computational resources, the work on generative models in the domain of remote sensing is still relatively scarce.
3.3 Capabilities
Inspired by MMBench [59], we have categorized the capabilities exhibited by VLGFMs into three hierarchical levels and clarified the design principles underlying this classification.
Figure 3: Overview of hierarchical capability levels of VLGFMs from $L_{0}$ to $L_{2}$. .
Overview. Figure 3 provides an overview of the existing multimodal capabilities of VLGFM. This benchmark encompasses 20 different leaf capabilities, offering a comprehensive evaluation of various fine-grained perceptual and reasoning abilities. Figure 4 illustrates each capability through specific examples, while Table V provides the capabilities supported by major remote sensing vision-language benchmarks. Since certain capabilities currently lack corresponding qualitative analysis datasets, they are not displayed in this table. Please note that the classification displayed is not the final version. We plan to expand the capability taxonomy further to make it more comprehensive.
Figure 4: Some examples of VLGFM’s capabilities.
The Hierarchical Taxonomy of Capability. Humans possess remarkable perceptual and reasoning abilities, enabling them to understand and interact with the world. These capabilities have been crucial in human evolution and are the foundation for complex cognitive processes. Perception involves gathering information from sensory inputs, while reasoning involves drawing conclusions based on this information. Collectively, they underpin the foundation of most tasks in the real world, including object recognition, problem-solving, and decision-making [84, 85]. VLGFMs are also expected to exhibit strong perceptual and reasoning capabilities in pursuit of artificial intelligence for remote sensing verticals. Therefore, our capability taxonomy designates perception and reasoning as our top-level capability dimensions, referred to as L-1 capability dimensions. For L-2 capabilities, we derive image-wise perception, region-wise perception, instance-wise perception, pixel-wise perception from L-1 perception, Attribute reasoning, and logical reasoning from L-1 reasoning. To make our benchmark as granular as possible, we further derive 20 L-3 capability dimensions from L-2.
3.3.1 Perception
Perceptual capability is the foundational skill for interpreting RSIs. With the rise of deep learning technology, numerous specialized models have been proposed to address image perception tasks in remote sensing, such as scene classification [38], object detection [13], and segmentation [86]. The emergence of VLGFM provides a solution for handling all perception tasks using a single model. Perception can be mainly divided into three types: image-wise, instance-wise, and pixel-wise perception.
Imgae-wise perception belongs to coarse-grained perception and is relatively easier. It can be divided into image scene, image quality, image modality, image captioning, and image retrieval. Instance-wise perception belongs to fine-grained perception, which requires identifying objects within an image and is more challenging. It can be divided into eight categories, encompassing some visual grounding capabilities. The basic building blocks of an image are pixels; thus, pixel-wise perception represents the most fine-grained perceptual capability. Pixel-wise perception often requires an additional pixel decoder. The SAM decoder [73] is a common choice. The output features of VLGFM will be aligned in dimensions and then concatenated with the prompt features of SAM. Finally, they are fed into the SAM decoder to generate the mask. Pixel-wise perception can be divided into referring expression segmentation and open-vocabulary segmentation.
$\bullet$ Image Scene aims to categorize an image into a predefined set of scenes or environments based on its overall visual content. This review, referencing the SATIN framework [38], categorizes remote sensing scenes into six distinct categories. Land cover refers to the biophysical surface characteristics of the Earth, with examples including forests, grasslands, and buildings. Land use describes the economic and social functions of areas, such as residential zones, industrial sectors, and parking lots, including bodies of water utilized for specific purposes. Hierarchical land use takes this a step further by classifying land use into more detailed levels of granularity, for instance, categorizing an area broadly as a transportation area, then more specifically as highway land, and even more precisely as a bridge. Complex scenes are images that contain one or more land use labels, indicating a blend of different uses or covers in one scene. Rare scenes represent unusual or infrequent categories that do not typically fall under standard land cover or land use definitions, such as hurricane damage or wind turbines. Lastly, false color scenes are images represented in false color, focusing on a single land use or land cover class, aiding in the distinct visualization of specific area characteristics.
$\bullet$ Image Quality aims to evaluate the condition of RSIs, taking into account factors such as imaging quality, noise levels [87], cloud cover [88], and low-light conditions [89]. This capability helps users filter out low-quality RSIs to prevent noise from contaminating the results. To date, no remote sensing benchmarks are specifically designed to test image quality capabilities.
$\bullet$ Image Modality aims to classify the modality of remote sensing imagery, such as optical, hyperspectral, and Synthetic Aperture Radar (SAR). This capability enables agent models to identify the modality of an image, thereby invoking the appropriate interpretation tools for that modality. Currently, no remote sensing benchmarks are specifically designed to test image modality capabilities.
$\bullet$ Image Captioning aims to produce a coherent and contextually relevant narrative in natural language that encapsulates the overall scene depicted in an image. Image captioning can further be categorized based on the input type into change captioning [90], multi-temporal image captioning, and video captioning [91]. The input for the change captioning task consists of two images. The objective is to summarize and describe the changes observed in the later image relative to the earlier one in the temporal sequence. Compared to change captioning, which involves two images, multi-temporal image captioning deals with a larger set of images and aims to encapsulate the progression of changes exhibited across a series of temporal images. In the case of video captioning, the input is a video sequence, and the task is to describe the events occurring within the video.
$\bullet$ Image Retrieval amis to find and retrieve images from a large remote sensing database that are similar or relevant to a given query image or text. Since this paper is written for vision-language models, it only involves image retrieval using text queries. In practical applications, users can quickly pinpoint specific RSIs through textual descriptions.
$\bullet$ Change Detection aims to identify areas or objects that have changed in RSIs of the same place at different times. Compared to change captioning capabilities, it focuses more on specific instances in the image than the entire image. Observing changes in specific targets is a more practical function in real-world applications. Levir-CC [92] and CDVQA [93] have analyzed changes in vegetation and buildings, which are used to track the progression of urbanization.
$\bullet$ Object Counting aims to identify and count the instances of specific objects within an image. Due to the high density of targets typically present in RSIs, manual techniques are time-consuming and labour-intensive. This makes counting the number of targets in RSIs a particularly valuable capability. For example, we can count animals, trees, buildings, and vehicles in RSIs [94].
$\bullet$ Object Presence aims to determine whether a specific target type exists within an image. Essentially, this capability requires the AI to answer whether the number of such targets is 0, which overlaps somewhat with the object counting capability. However, the ability to detect the presence of targets is a common requirement in real-world applications, which is why it is recognized as a distinct leaf capability.
$\bullet$ Spatial Relationship aims to describe the absolute positional relationships of targets relative to the image or their relative positional relationships to other targets within the image. In the field of general vision-language foundation models, this capability is often considered a form of reasoning ability. However, because the positional relationships of targets in RSIs are relatively simpler, we categorize it as a perceptual capability.
$\bullet$ Attribute Recognition aims to identify the attributes of the specific target within an image, including various aspects of its appearance, such as colour, size, and material. For instance, detailed attribute features of a car can help determine the vehicle’s brand and even its specific model. This capability is highly dependent on the resolution of RSIs; higher resolution allows for capturing more detailed attributes of the target.
$\bullet$ Open-vocabulary Detection aims to detect and recognize objects in images, even those not encountered during training, by leveraging a flexible and expansive vocabulary. The open-vocabulary detection task can be considered an advanced version of the traditional object detection task. It detects common object categories in remote sensing object detection datasets and identifies objects with certain descriptive constraints. For instance, one can detect all vehicles in an image or, specifically, all white vehicles in an image. The detection results are returned in textual format.
$\bullet$ Referring Expression Comprehension aims to localize a target object in an image described by a referring expression phrased in natural language. The difference from open-vocabulary detection is that the referring expression comprehension refers to a single target, whereas the output of open-vocabulary detection can include multiple targets of the same type. The detection results are returned in the same text format as the training data.
$\bullet$ Referring Expression Generation aims at generating a discriminative description of an object in an image, which is very similar to the image captioning task. The distinction from attribute recognition lies in the fact that attribute recognition involves describing the features of a single target, whereas referring expression generation involves describing the contents within a specified bounding box.
$\bullet$ Referring Expression Segmentation aims to segment a target object in an image described by a natural language referring expression. It can be considered as the semantic segmentation version of Referring Expression Comprehension. Unlike Referring Expression Comprehension, its output is a binary mask image.
$\bullet$ Open-vocabulary Segmentation aims to segment and classify image regions using a comprehensive and extendable set of labels not limited by a predefined vocabulary. Unlike Open-vocabulary detection, its output is a binary mask image. RSIs encompass visible light and multispectral and hyperspectral data. Effectively utilizing this information for segmentation presents a significant challenge.
3.3.2 Reasoning
However, perception tasks can be accomplished using pure vision specifical models, while autonomous reasoning is the capability most eagerly anticipated in the field of remote sensing interpretation [95]. Reasoning poses greater difficulty and has a higher application value than perception. Reasoning can be mainly divided into two types: logic and attribute reasoning.
Logical reasoning requires leveraging expert knowledge from specific domains to complete the inference process. This external information can include geographical knowledge, cartographic knowledge, anthropological knowledge, agricultural knowledge, and more. Logical reasoning capabilities fully demonstrate the advantages of VLGFMs. Logical reasoning can be divided into geospatial localization, geometric measurement, and resource assessment. Compared to Attribute Recognition in perceptual capabilities, the attributes in Attribute Reasoning cannot be directly observed from the image but require inference combined with sensor parameters and physical knowledge. Attribute reasoning can be divided into Sensor Property and Physical Property.
$\bullet$ Geospatial Localization aims to pinpoint the geographic location represented in a remote sensing image, potentially identifying the specific continent, country, city, or precise latitude and longitude based on terrain and landmark features [96]. Its inspiration comes from the popular geography game GeoGuessr [^4] https://www.geoguessr.com, in which players are placed in a street view panorama somewhere in the world and must guess their location. Geospatial Localization typically necessitates acquiring information on topography, vegetation, and building styles through perceptual capabilities to perform reasoning, which adds significant complexity.
$\bullet$ Geometric Measurement aims to measure the physical dimensions of objects and features on the ground using RSIs, including estimating height, distance, and calculating area. The application of this capability is extensive; for example, it can be used to estimate the distance between two buildings in an image, the relative elevation from the foot to the summit of a mountain, the length of the ship, and the area occupied by cropland [97].
$\bullet$ Resource Assessment aims to estimate and analyze natural resource quantity, quality, and spatial distribution, such as estimating urban population sizes [95], calculating forest carbon dioxide emissions, and estimating crop yields.
$\bullet$ Sensor Property aims to infer the parameters of the imaging sensor, which includes documenting the time of capture, the altitude and direction of the sensor when the image was taken, and the camera’s tilt angle and focal length. Due to the non-intuitive impact of different remote sensing sensor parameters, inferring sensor properties from images is not straightforward. However, the relevant parameters of the sensor are crucial for tasks such as oblique photography and monocular 3D reconstruction.
$\bullet$ Physical Property aims to reason the physical characteristics or conditions of objects within an image, such as the moving targets or the houses submerged by a flood. The physical properties of the target are often not directly obtainable from the image and require inference combined with physical common sense. For example, if rust is observed on the roof of a building in a remote sensing image, the roof is likely made of iron. If a ship is sailing in the water, it will leave a wake behind it.
TABLE V: Capabilities supported by major remote sensing vision-language benchmarks. Abbreviations IS, IC, IR, CD, OP, OC, SR, AR, REC, REG, RES, OVS, GM and PP in the table stand for image scene, image captioning, image retrieval, change detection, object presence, object counting, spatial relationship, attribute recognition, open-vocabulary detection, referring expression comprehension, referring expression generation, referring expression segmentation, open-vocabulary segmentation, geospatial localization, geometric measurement, resource assessment, sensor property, and physical property, respectively.
Dataset IS IC IR CD OP OC SR AR REC REG RES OVS GM PP Total
Sydney-Captions [40] 1
UCM-Captions [40] 2
RSICD [43] 2
RSVQA-LR [41] 5
RSVQA-HR [41] 6
RSVQAxBEN [98] 2
FloodNet [99] 2
RSITMD [39] 2
RSIVQA [100] 4
NWPU-Captions [44] 1
LEVIR-CC [92] 2
CDVQA [93] 4
UAV [46] 1
RSVG [42] 5
CRSVQA [101] 5
DIOR-RSVG [47] 4
RSIEval [24] 6
SATIN [38] 1
NAIP-OSM [26] 2
RemoteCount [22] 1
SkyScript [27] 2
EarthVQA [102] 5
RRSIS [103] 1
GeoChat-Bench [25] 6
RRSIS-D [104] 1
VLEO-Bench [94] 6
FineGrip [105] 2
4 Challenges and Outlook
4.1 Challenges
$\bullet$ Image resolution too low. The spatial resolution of RSIs determines the actual ground area represented by each pixel. A low resolution implies that each pixel covers a larger ground area. Due to the limitations of current graphics card performances and the extensive parameter size of large visual-language models, it is not yet feasible to input high-resolution RSIs without crop. In remote sensing imagery, identifying significant geographical and environmental features, such as vehicles, trees, and types of crops, necessitates high resolution. These details may appear blurred in low-resolution images, making it difficult for models to recognize ground objects accurately. In remote sensing visual-language tasks, the model needs to describe the content of the images precisely. Insufficient information provided by low-resolution images may result in inaccurate or detail-lacking language descriptions, adversely affecting the overall performance and credibility of the model.
$\bullet$ Training cost too high. As shown in Table I, VLGFMs have high hardware resource requirements, with most work completed using A100 GPUs. However, for some university laboratories, the cost of A100 GPUs is prohibitively high. LMDeploy [106] is attempting to reduce memory requirements during inference through quantization. However, the high demand for GPUs during the training phase inevitably raises the barrier for research in this direction.
$\bullet$ Benchmark too limited. On one hand, as shown in Table V, the existing benchmarks do not comprehensively cover various capabilities, making it difficult to measure the performance of VLGFMs fully. On the other hand, the tasks in existing multimodal benchmarks are relatively simple, with most humans being able to solve them easily. Hence, even if VLGFMs achieve high scores on these benchmarks, it is still not convincing.
$\bullet$ Correction of hallucinations. LLMs tend to generate absurd or untrue content unrelated to any sources, a phenomenon researchers call hallucination [107]. Due to the presence of hallucinations, current large models are still not widely applicable in high-precision tasks such as critical target interpretation and remote sensing mapping. To mitigate the hallucination phenomenon, researchers train models to refuse to answer questions they cannot handle. However, some researchers have a different perspective on the hallucination phenomenon. They argue that hallucination is not only a problem but also a potential capability. This capability can help LLMs generate unexpectedly interesting answers. Whether considered a capability or a challenge, the hallucination phenomenon in LLMs is an issue that must be addressed as we work to enhance model capabilities and practical applications.
4.2 Future Work
$\bullet$ More powerful LLMs. In visual-language models, the language model is responsible for parsing and generating linguistic descriptions, which is crucial for understanding the content and context of images. If the language model can accurately and deeply understand and generate natural language, it will more effectively integrate with visual data, providing richer and more accurate interpretations of information. Language models typically contain a vast amount of world knowledge and common sense, which are vital for interpreting complex visual scenes. Lastly, if there are biases or misunderstandings in the language model, these errors could be amplified in the visual-language model. Therefore, the language model’s accuracy directly affects the visual-language model’s reliability and effectiveness. Excellent language understanding and generation capabilities are fundamental to ensuring the success of visual-language models. Recently, GPT-4o [^5] https://openai.com has been released by OpenAI. It not only achieves native multimodal capabilities but also supports end-to-end speech input and output. Compared to agent-based methods, it offers lower latency and an improved user experience. This signifies an advancement in LLMs’ interactive capabilities, opening up more possibilities for the future development of VLGFM.
$\bullet$ More challenging benchmarks. When models perform well on simple tasks, their underlying problems and limitations may not be apparent. More challenging benchmarks can test the performance of models under extreme or unknown conditions, thus helping researchers identify and address defects within the models, enhancing their robustness and generality. Real-world application scenarios are often more complex and varied than standard datasets. By testing and optimizing visual-language models on more challenging benchmarks, models can be better prepared to meet the demands of the real world, improving their practicality and reliability.
$\bullet$ Training-free technology. Given that the high training costs raise the barrier for VLGFMs, achieving good results in the remote sensing without extensive training would be a welcome development for researchers. This approach is referred to as training-free technology. Chain of thought (CoT) [108] might be one of the more promising training-free approaches. In this method, the model generates a reasoning process before providing an answer, similar to how humans often make correct judgments after careful consideration. Zero-shot CoT [109] employs a single, task-agnostic ”general” trigger to prompt the model to generate a reasoning chain. In the first stage, a phrase such as ”Let’s think step by step” is appended to the question and input into the LLM, prompting it to generate a step-by-step logical reasoning process for the problem. Next, the prompt from the first stage and the reasoning chain generated by the model are concatenated and used as the input for the second stage, where the model is then asked to generate the answer. However, this capability to generate logical chains only emerges when the model reaches a certain scale. It is worth mentioning that the aforementioned methods do not require training, which significantly lowers the research barrier.
$\bullet$ More powerful interpretability. In the field of remote sensing intelligence interpretation, interpretability is of paramount importance. It enables researchers to delve deeply into the behavior of models, thereby identifying and correcting errors or biases therein. In remote sensing data analysis, erroneous predictions can stem from various factors such as data quality, model architecture, or insufficient training. By interpreting the outputs of models, researchers can more effectively adjust and optimize these models. Additionally, the interpretation of remote sensing data often requires expertise from the fields of geographic science, ecology, and other disciplines. An interpretable model aids experts from diverse backgrounds in better understanding the outputs of the model, facilitating interdisciplinary communication and collaboration, thereby enhancing the accuracy and practicality of the analysis. In summary, interpretability enhances the usability and reliability of remote sensing intelligence interpretation systems.
5 Summary and conclusions
This survey provides a detailed examination of the latest developments in VLGFMs, which appear to be the first of its kind. We outline the necessary background knowledge, including the fundamental concepts and introductory information about VLGFM. Subsequently, we summarize two data collection methods, three model architectures, and twenty fundamental capabilities. For each capability, we provide straightforward descriptions and intuitive examples. Finally, we summarize several challenges and highlight some future research directions for VLGFMs.